The maximum amount of silver carbonate that will dissolve in a 0.169 M potassium carbonate solution is
Answer #1

Answer

molar solubility of Ag2CO3 in 0.169M K2CO3=3.47×10-6 mol/L

solubility of Ag2CO3 in 0.169M K2CO3 = 9.57×10-4g/L

Explanation

Solubility equillibrium of Ag2CO3 is

Ag2CO3(s) <-------> 2Ag+(aq) + CO32-(aq)

Ksp = [Ag+]2[CO32-] = 8.10 ×10-12

substitute the given concentration of CO32- is Ksp -expression

[Ag+]2 × 0.169M = 8.12 ×10-12M3

[Ag+]2 = 4.805 ×10-11M2

[Ag+] = 6.932 ×10-6M

molar solubility of Ag2CO3 = [Ag+] /2 = 6.932×10-6M/2= 3.47×10-6 mol/L

mass of 3.47 ×10-6moles of Ag2CO3 = 3.47 ×10-6mol × 275.75g/mol = 9.57×10×10-4g

solubility of Ag2CO3 = 9.57 ×10-4g/L